OpenStudy (trojanpoem):

Using Wronskian determinant: Check the following for linear dependency:

OpenStudy (trojanpoem):

1. e^-x, e^x , e^2x

OpenStudy (trojanpoem):

Apparently , they are not dependent , but how to prove that with Wronskain ?

OpenStudy (phi):

did you write down the wronskian ? what does it look like?

OpenStudy (phi):

W has to be square you need 3 rows f(x) f'(x) f''(x) where f(x) is the row of initial functions then evaluate it at e.g. x=0

OpenStudy (trojanpoem):

I don't know how to write it ... All I know is for something like x^2, x^3 g(x) = x^2 g'(x) = 2x f(x) = x^3 f'(x) = 3x^2

OpenStudy (trojanpoem):



OpenStudy (trojanpoem):

This one is a determinant because it's a square matrix.

OpenStudy (trojanpoem):

But in the given case ? I have matrix 2 x 3

OpenStudy (phi):

in this case 1. e^-x, e^x , e^2x the 2nd row is the first derivative the 3rd row is the 2nd derivative

OpenStudy (trojanpoem):



OpenStudy (phi):

yes, now evaluate at x=0 (easy value to use) 1 1 1 -1 1 2 1 1 4

OpenStudy (trojanpoem):

1 * ( 4 - 2) - 1 * ( -4 - 2) + 1 * ( -1 - 1 ) 2 -1 * -6 -2 2 +6 - 2 = 6

OpenStudy (trojanpoem):

not zero, so they are not dependent.

OpenStudy (trojanpoem):

But phi ? can I use any value ?

OpenStudy (phi):

to be dependent , W must be 0 for all values if you find just one counter-example, they are not dependent.

OpenStudy (phi):

*det(W) not just W

OpenStudy (trojanpoem):

How can I check for all values ?!

OpenStudy (phi):

You don't. You find one counter example, thus showing the functions are not dependent, and therefore they are independent.

OpenStudy (trojanpoem):

Thanks Phi.